OpenStudy (anonymous):

Please help? I don't understand, will give medal. The following function defines a recursive sequence: f(0) = -3 f(1) = 6 f(n) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? A. -3, 6, -9, -12, … B. -3, 20, -95, 480, … C. -3, 6, -9, 12, … D. -3, -20, -95, -480, … I believe it is either A, or C.

OpenStudy (anonymous):

\[f(n) = -2f(n -1) - f(n - 2)\]

OpenStudy (anonymous):

this means the next term is two times the previous term minus the term before that

OpenStudy (anonymous):

ok, f(0) = -3 , first term = -3 f(1) = 6, 2nd term = 6 so, as of now our sequence is -3,6 to find nect term, we will use f(n) = -2•f(n -1) - f(n - 2) put n= 2 here, what do u get ?

OpenStudy (anonymous):

you just replace "n" by "2" there, to get f(2)

OpenStudy (anonymous):

f(2) = -2 f(1) - f(0) and you already know f(1) =6, f(0) = -3 just plug in values!

OpenStudy (anonymous):

you need to find f(3) too! f(3) = - 2f(2) - f(1) = - 2(-9) -6 = 18-6 = 12

OpenStudy (anonymous):

oops \[f(2)=-2f(1)-f(0)=-2\times 6-(-3)\]

OpenStudy (anonymous):

So then it would be C ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah it is always C

OpenStudy (anonymous):

Thank you so much! Both of you, wish I could give you both medals :/

OpenStudy (anonymous):

who help you the most is the one you should give a medal too

OpenStudy (anonymous):

if you felt @satellite73 help u the most give it to satellite

OpenStudy (anonymous):

lol i have plenty, no worries

OpenStudy (anonymous):

okay lol

OpenStudy (anonymous):

plus @lupita1995 is most helpful here

OpenStudy (anonymous):

awwwwwwwwwwwwwwwww

OpenStudy (anonymous):

thank you